# Spatially–Structured Models of Viral Dynamics: A Scoping Review

### v1

@@ 4.4 Agent–based modelling: cellular automata
Figure 8: Cellular automata summary. Seminal publications are as follows: Bernaschi and Castiglione [32], Zorzenon dos Santos and Coutinho [224], Beauchemin et al. [27], Xiao et al. [213], Heilmann et al. (2010) [93], Wodarz et al. [210], and Heilmann et al. (2012) [94].
@@ 4.5 Agent–based modelling: multicellular models
Figure 9: Multicellular models summary. Seminal publications are as follows: Mitchell et al. [145], Yakimovich et al. [215], Eriksen et al. [66], Imle et al. [99], Getz et al. [78], Simmons et al. [190], and Sego et al. [183].
@@
The simplest approach under the multicellular framework is in coupling a cellular automaton to a global virus population. This construction has been used by authors in a variety of applications (Cangelosi et al. [45] and Goyal and Murray [82] in HBV infection, Blahut et al. [37] in HCV, Kreger et al. [121] and Shi et al. [188] in HIV, for example) and generally follows from an assumption that free viruses spread quickly relative to the size of the model tissue [45, 82, 37, 121]. A similar idea was applied by Howat et al., who considered the spread of free virus at the cell scale, but type I interferon (comprised of molecules far smaller than virions which therefore diffuse much more quickly) as a global quantity [97]. Models which consider only infection from a global viral reservoir do not give rise to spatially–structured infections, and as such, the global viral population is incorporated into the model dynamics in different ways. Shi et al. considered neighbour–based infection at a rate related to the global viral load [188]. More typically, however, models with a global virus population generally consider two modes of viral infection: unstructured infection due to the global population of free virions, and structured local infection arising from direct cell–to–cell infection between neighbouring cells (Saeki and Sasaki [173], Kreger et al. [121], Goyal and Murray [82], Blahut et al. [37]). Such models therefore describe two spatial scales of infection, which are generally not well–described by cellular automata. We describe modelling of the cell–to–cell infection route in further detail in Section 6.
@@ 5.1 HIV
Figure 10: Spatially–structured models for HIV. Seminal publications are as follows: Zorzenon dos Santos and Coutinho [224], and Imle et al. [99].
@@ 5.2 Oncolytic viruses
Figure 11: Spatially–structured models for oncolytic viruses. Seminal publications are as follows: Wu et al. [212], Wein et al. [205], Friedman et al. [73], Mok et al. [146], Wodarz et al. [210], Malinzi et al. [132], Friedman and Lai [72], and Alzahrani et al. [9].
@@
A popular application of models for oncolytic viral infections is in testing the impact of different approaches to viral administration to the tumour. Wu et al. [212] and Wein et al. [205] both studied models for radially symmetric tumours with virus initially delivered either to the rim or to the core of the tumour. Wu et al. suggested that, taking into account the necrosis of the tumour core due to nutrient diffusion from the rim, core injection was likely to be somewhat redundant, and that rim injection would be more effective in spreading through the tumour [212]. Wein et al. found that for oncolytic virotherapy to be effective, the virus must be highly aggressive — that is, extremely infectious and fast–spreading — in order to outpace the growth of the tumour [205]; this result was recapitulated by Al-Johani et al. [4] and Simbawa et al. [189]. Pooladvand et al., however, has warned that high infectivity is not necessarily sufficient for complete eradication of the tumour [164]. Other authors have also found that the success of oncolytic virotherapy depends on viral infection reaching widespread regions of the tumour. Storey and Jackson compared a viral dosage strategy which repeatedly targeted the same location in the tumour to an adaptive strategy which instead delivered oncolytic viral load to the location of highest tumour cell density at that point in time, and found that the adaptive strategy dramatically reduced the final tumour mass after the same length of treatment time [195]. Bailey et al. proposed a model with randomly seeded foci arising from viral extravasation but with a hard limit on the radius of infection foci, following experimental observations of limited focal spread [20]. This approach suggested that, in order for virotherapy to be effective, the number of viral foci would have to be extremely high, such that the overall infected volume would cover a sufficiently wide region [20].
@@
This spatially constrained spread of infection foci within tumours is a recurring theme in explaining the limited in vivo efficacy of oncolytic viruses. One reason for this limitation in the spatial spread of infection is due to the fact that the extremely dense tumour environment severely inhibits the diffusion of free virions [110, 147]. As a result, many authors have argued that infection is likely only possible directly between neighbouring cells (Morselli et al. [147], Bhatt et al. [36], Reis et al. [168]). Wodarz et al. showed that experimentally–observed infection patterns for oncolytic viruses could be generated from a cellular automaton with adjacency–based infection only [210]. As a result, it has been widely argued that increased diffusion of free virions would dramatically improve oncolytic virotherapy (Paiva et al. [155], Jenner et al. [110, 109], Mok et al. [146], for example), however, there is disagreement in the modelling community about the conditions under which this might come about. Mok et al. [146] and Paiva et al. [155] have argued that a reduction in density of the tumour extracellular matrix would increase the rate of viral diffusion, however, Alsisi et al. [6, 7, 5] have suggested that virions preferentially migrate towards regions of higher matrix density via haptotaxis. A related problem to the rate of viral diffusion is the relative density of the tumour cells and normal stromal cells. Since the latter cannot be infected by the oncolytic virus, Wein et al. [205] and Jenner et al. [110] have shown that oncolytic viral infections are far more effective in tumours where cancer cells are densely packed, with few interspersed stromal cells. However, in contrast to these findings, Berg et al. showed that oncolytic infections spread more slowly in a three–dimensional tumour model compared to an equivalent model in two dimensions, despite cells in the former having a higher number of adjacent tumour cells [31].
@@
Since the spread of viral infections in tumours is mainly between neighbouring cells, the arrangement of cells in space is highly important in agent–based modelling of oncolytic viruses. Simpler approaches to the arrangement of tumour cells in space consider a regular lattice of cell–sites which remain fixed over time (Wodarz et al. [210], Reis et al. [168], Paiva et al. [155] for example). Others have constructed spatially–static networks of cells based on Voronoi tessellations derived from experimental imaging of tumours, which provides a realistic distribution of the number of neighbours between cells across the tumour (Berg et al. [31] and Bhatt et al. [36]). Berg et al. compared oncolytic viral dynamics on a regular and Voronoi cell lattice in three dimensions and found that infections on the Voronoi lattice were less likely to result in eradication of the tumour [31]. Temporal evolution of the cell grid has also been implemented. Jenner et al. considered a time–varying Voronoi lattice model for tumour cells, which also enabled the explicit representation of tumour cell proliferation and remodelling, and moreover facilitated the simulation of time–varying adjacency between cells [109].
@@ 5.3 Respiratory viruses
Figure 12: Spatially–structured models for influenza and SARS–CoV–2. Seminal publications are as follows: Beauchemin et al. [27], Mitchell et al. [145], Getz et al. [78], Quirouette et al. [166], and Sego et al. [183].
@@
Viruses which infect the respiratory tract specifically target the monolayer of epithelial cells which line the airways [53, 149, 76]. Thus, to an extent, the natural geometry of respiratory infection is two–dimensional, especially at the microscopic scale, and as a consequence virtually all of the respiratory infection models retained for analysis from our search also consider two spatial dimensions. However, at a larger length scale, there is substantial heterogeneity in the structure of the lung along the length of the respiratory tract. The respiratory tract has a complex, tree–like structure, comprising 24 generations of branching airways, such that upper passages like the trachea are relatively wide (on the order of centimetres), while deeper airways are very narrow (on the order of fractions of millimetres) [130, 53]. Ayadi et al. [16] and Moses et al. [149] have explicitly considered the dynamics of viral infection across the entire, heterogeneous lung. In the latter case, Moses et al. constructed a full–size, three–dimensional digital human lung, complete with a detailed branching structure for the cells [149]. The authors showed that viral dynamics on this highly complex structure were significantly different to the dynamics of the same infection model on a purely two–dimensional cell grid [149]. While the use of this hyperrealistic geometry demonstrates the notable differences in the dynamics of infections on flat sheets compared to the lung, there is a substantial conceptual leap between the two geometries. Other authors have instead employed an intermediate approach, studying the spread of viral infections on sections of the lung in order to identify specific ways in which the structure and geometry of lung tissue influence the spread of infection. Chen et al. [53] and Williams et al. [208] studied the spread of viral infections on individual lung branches, or on small sections of branching tissue. The latter group showed that tubular and branching tissue geometry imposes rich, spatially–directional infection dynamics [208]. These dynamics moreover exhibited an increased sensitivity to the rate of viral diffusion compared to the dynamics of infection on a flat cell sheet [208].
@@
Another characteristic feature of the lung is the mucociliary escalator, the directed flow of the mucus lining the respiratory epithelium from lower to upper regions of the respiratory tract. Quirouette et al. [166] included the mucociliary escalator in a model of influenza infection, and suggested that infection can effectively only spread up the respiratory tract from the initial point of virion deposition [166]. The rate of mucociliary flow is variable at different depths of the respiratory tract, with slower flow in deeper regions of the lung [130, 53]. Chen et al. compared the dynamics of SARS–CoV–2 infection in airway tissue from different generations of the respiratory tract and suggested that infection spread was severely limited in the lower airways [53]. The authors postulated that infection propagation in the deep lung must therefore be driven by a large quantity of viral depositions from upper regions of the lung, such as virus–laden droplets or aerosols [53]. Chakravarty et al. — who extended the Quirouette model — reached a similar conclusion, arguing that aerosolisation of infected nasopharyngeal mucosa (ANM) was necessary for infection to reach the deep lung [52]. Vimalajeewa et al., by contrast, modelled advection down the respiratory tract in the course of SARS–CoV–2 infection due to air flow in the respiratory lumen (the authors accounted for the mucociliary escalator as a drag force and slip correction factor in the flow model) [203]. In addition to its role in disease progression, Zhang et al. suggested that the rate of mucus advection also influences the nasal viral load [221]. The authors argue that variation in this rate between individuals could offer one potential explanation for the extreme variation observed in clinical nasal swab titres among COVID–19 patients [221].
@@ 5.4 Other viruses
Figure 13: Spatially–structured models for other specific viruses. Seminal publications are as follows: Xiao et al. [213], Heilmann et al. (2010) [93], Heilmann et al. (2012) [94], Schiffer et al. [177], Eriksen et al. [66], and Simmons et al. [190].
@@
In addition to the viruses which we have discussed above, several other species have been considered. Among these, a substantive body of work has been developed for the modelling of bacteriophage (or simply “phage”) dynamics. Phage and the bacteria which they infect are both some of the largest biological populations in existence, and their interaction is one of the oldest predator–prey relationships on Earth [190, 144]. Bacteria and phage inhabit and interact within a broad array of environments, including in soil, water, and within larger organisms [93, 77]. We note that there is a major difference between phage dynamics and the dynamics of the other virus species which we have considered in this review, in that phage hosts are single–cell organisms. As such, while models for phage–bacteria interactions describe the spread of infection among cells — as with the other models we have discussed — such systems are, strictly, between–host models. We have nonetheless included such studies in our analysis.
@@
Due to the unique properties of phage–bacteria dynamics, as compared to the dynamics of other viral species we have discussed, modellers have considered the spatially–structured interactions between phage and microbe communities on far larger spatiotemporal scales. Garrido Zornoza et al. modelled the spread of phage among a microbial community at an ecological spatial scale (the only such model we found in our search), which described spatially explicit microbial habitats, from which both viruses and host microbes were subject to stochastic aerosolisation and advection with wind [77]. The authors showed that this influx–outflux process among habitats lead to the spatial spread of microbial communities but also resulted in destabilisation and occasional local extinctions in individual habitats [77]. While phage–bacteria systems exhibit spatially–structured dynamics on this ecological spatial scale, most authors instead model the dynamics of phage–bacteria interactions on far smaller length scales. In particular, many authors have modelled the spread of phage infection among dense, surface–bound microbial communities called biofilms. The crowded nature of these environments requires particular consideration by modellers. Martínez-Calvo et al., for instance, considered density-dependent motility of bacteria and showed that the spread of phage infection among the colony can disrupt bacterial aggregation [138]. Moreover, as with oncolytic viral infections in tumours, the densely–packed environment of biofilms almost entirely suppress viral diffusion, resulting in highly localised viral spread. Taylor et al., for example, showed that the high cell density found in biofilms promotes the formation of spatial clustering of infection [198]. Modellers have accounted for the crowded nature of biofilms in various ways. While some, including Heilmann et al. (2012) [94] and Mitarai et al. [144], have implemented linear diffusion to describe the dispersal of cell–free phage with a very small diffusion coefficient compared to say, the Stokes–Einstein estimate (Akpinar et al. [3], Beauchemin et al. [26], Sego et al. [183], for example, took similar approaches for other viral species). Other authors have instead opted for non–linear forms of the diffusion term: Hunter et al., for example, considered density–dependent diffusion [98]; Simmons et al. explicitly considered physical blocking of phage motion by bacteria [190]. Valdez et al. showed that, as a consequence of this highly localised spread, differently–shaped initial distributions of phage among bacterial communities gave rise to spatially distinct colony structures [202].
@@
Hepatitis B and C viruses, which cause chronic infections of the liver, have also been modelled by several authors. Despite the recent development of highly effective direct–acting antivirals for the latter, hepatitis viruses are still globally widespread with relatively high rates of mortality [90, 58]. Xiao et al. [213] and Cangelosi et al. [45] have argued that spatial heterogeneity of infection may play a role in immune evasion, supporting viral persistence in the liver. Goyal and Murray [82] and Blahut et al. [37] have also argued that direct cell–to–cell transmission of hepatitis viruses also contributes to the chronicity of infection (we discuss modelling for cell–to–cell infection in greater depth in Section 6). Moreover, within–cell dynamics appears to be key in hepatitis B dynamics. Cangelosi et al. showed that viral cccDNA represents one of the most robust obstacles to hepatitis B treatment, and suggested that even a single cccDNA molecule would be sufficient to reinfect the entire liver [45].

### v2

@@ 4.4 Agent–based modelling: cellular automata
Figure 8: Cellular automata summary. Left: Average number of publications using cellular automata per year. Centre: Cellular automaton publications by viral species studied. Right: Highly–cited cellular automaton publications over time. Highly–cited publications are as follows: Bernaschi and Castiglione [24], Zorzenon dos Santos and Coutinho [25], Beauchemin et al. [29], Xiao et al. [31], Heilmann et al. (2010) [33], Wodarz et al. [35], and Heilmann et al. (2012) [36].
@@ 4.5 Agent–based modelling: multicellular models
Figure 9: Multicellular models summary. Left: Average number of publications using multicellular models per year. Centre: Multicellular models by viral species studied. Right: Highly–cited multicellular model publications over time. Highly–cited publications are as follows: Mitchell et al. [34], Yakimovich et al. [5], Eriksen et al. [41], Imle et al. [2], Getz et al. [9], Simmons et al. [44], and Sego et al. [8].
@@
The simplest approach under the multicellular framework is in coupling a cellular automaton to a global virus population. This construction has been used by authors in a variety of applications (Cangelosi et al. [153] and Goyal and Murray [99] in HBV infection, Blahut et al. [107] in HCV, Kreger et al. [4] and Shi et al. [154] in HIV, for example) and generally follows from an assumption that free viruses spread quickly relative to the size of the model tissue [153, 99, 107, 4]. A similar idea was applied by Howat et al., who considered the spread of free virus at the cell scale, but type I interferon (comprised of molecules far smaller than virions which therefore diffuse much more quickly) as a global quantity [155]. Similarly, Durso-Cain et al. considered a diffusing viral population but a global antibody population [156]. Models which consider only infection from a global viral reservoir do not give rise to spatially–structured infections, and as such, the global viral population is incorporated into the model dynamics in different ways. Shi et al. considered neighbour–based infection at a rate related to the global viral load [154]. More typically, however, models with a global virus population generally consider two modes of viral infection: unstructured infection due to the global population of free virions, and structured local infection arising from direct cell–to–cell infection between neighbouring cells (Saeki and Sasaki [3], Kreger et al. [4], Goyal and Murray [99], Blahut et al. [107]). Such models therefore describe two spatial scales of infection, which are generally not well–described by cellular automata. We describe modelling of the cell–to–cell infection route in further detail in Section 6.
@@ 5.1 HIV
Figure 10: Spatially–structured models for HIV. Left: Average number of models for HIV per year. Centre: HIV models by model type. Right: Highly–cited publications for HIV models over time. Highly–cited publications are as follows: Zorzenon dos Santos and Coutinho [25], and Imle et al. [2].
@@ 5.2 Oncolytic viruses
Figure 11: Spatially–structured models for oncolytic viruses. Left: Average number of models for oncolytic viruses per year. Centre: Oncolytic virus models by model type. Highly–cited publications for oncolytic virus models over time. Highly–cited publications are as follows: Wu et al. [26], Wein et al. [27], Friedman et al. [30], Mok et al. [32], Wodarz et al. [35], Malinzi et al. [39], Friedman and Lai [40], and Alzahrani et al. [42].
@@
A popular application of models for oncolytic viral infections is in testing the impact of different approaches to viral administration to the tumour. Wu et al. [26] and Wein et al. [27] both studied models for radially symmetric tumours with virus initially delivered either to the rim or to the core of the tumour. Wu et al. suggested that, taking into account the necrosis of the tumour core due to nutrient diffusion from the rim, core injection was likely to be somewhat redundant, and that rim injection would be more effective in spreading through the tumour [26]. Wein et al. found that for oncolytic virotherapy to be effective, the virus must be highly aggressive — that is, extremely infectious and fast–spreading — in order to outpace the growth of the tumour [27]; this result was recapitulated by Al-Johani et al. [184] and Simbawa et al. [185]. Pooladvand et al., however, has warned that high infectivity is not necessarily sufficient for complete eradication of the tumour [186]. Other authors have also found that the success of oncolytic virotherapy depends on viral infection reaching widespread regions of the tumour. Storey and Jackson compared a viral dosage strategy which repeatedly targeted the same location in the tumour to an adaptive strategy which instead delivered oncolytic viral load to the location of highest tumour cell density at that point in time, and found that the adaptive strategy dramatically reduced the final tumour mass after the same length of treatment time [187]. Bailey et al. proposed a model with randomly seeded foci arising from viral extravasation but with a hard limit on the radius of infection foci, following experimental observations of limited focal spread [67]. This approach suggested that, in order for virotherapy to be effective, the number of viral foci would have to be extremely high, such that the overall infected volume would cover a sufficiently wide region [67]. However, since the model is phenomenological, the authors cannot deduce a mechanistic explanation for the failure of the treatment at a reasonable dosage.
@@
This spatially constrained spread of infection foci within tumours is a recurring theme in explaining the limited in vivo efficacy of oncolytic viruses. One reason for this limitation in the spatial spread of infection is due to the fact that the extremely dense tumour environment severely inhibits the diffusion of free virions [170, 145]. As a result, many authors have argued that infection is likely only possible directly between neighbouring cells (Morselli et al. [145], Bhatt et al. [180], Reis et al. [141]). Wodarz et al. showed that experimentally–observed infection patterns for oncolytic viruses could be generated from a cellular automaton with adjacency–based infection only [35]. As a result, it has been widely argued that increased diffusion of free virions would dramatically improve oncolytic virotherapy (Paiva et al. [151], Jenner et al. [170, 165], Mok et al. [32], for example), however, there is disagreement in the modelling community about the conditions under which this might come about. Mok et al. [32] and Paiva et al. [151] have argued that a reduction in density of the tumour extracellular matrix would increase the rate of viral diffusion, however, Alsisi et al. [72, 73, 74] have suggested that virions preferentially migrate towards regions of higher matrix density via haptotaxis. A related problem to the rate of viral diffusion is the relative density of the tumour cells and normal stromal cells. Since the latter cannot be infected by the oncolytic virus, Wein et al. [27] and Jenner et al. [170] have shown that oncolytic viral infections are far more effective in tumours where cancer cells are densely packed, with few interspersed stromal cells. However, in contrast to these findings, Berg et al. showed that oncolytic infections spread more slowly in a three–dimensional tumour model compared to an equivalent model in two dimensions, despite cells in the former having a higher number of adjacent tumour cells [188]. The lack of consensus in these findings points to complex, multifactorial dynamics at play, and suggests a need for an improved understanding of the various mechanisms of oncolytic virus dispersal within tumours, including through cell–to–cell transmission, diffusion, and haptotaxis. Moreover, these results suggest that a realistic representation of the tumour microenvironment, such as the extracellular matrix density, the distribution of cancer cells and non–cancer cells, and perhaps the dimensionality of the tumour, may be necessary to ensure dynamics faithful to oncolytic viral infection in vivo.
@@
Since the spread of viral infections in tumours is believed to mainly between neighbouring cells, the arrangement of cells in space is likely highly important in agent–based modelling of oncolytic viruses. Simpler approaches to the arrangement of tumour cells in space consider a regular lattice of cell–sites which remain fixed over time (Wodarz et al. [35], Reis et al. [141], Paiva et al. [151] for example). Others have constructed spatially–static networks of cells based on Voronoi tessellations derived from experimental imaging of tumours, which provides a realistic distribution of the number of neighbours between cells across the tumour (Berg et al. [188] and Bhatt et al. [180]). Berg et al. compared oncolytic viral dynamics on a regular and Voronoi cell lattice in three dimensions and found that infections on the Voronoi lattice were less likely to result in eradication of the tumour [188]. Spatiotemporal evolution of the cell grid is also likely important in the dynamics of oncolytic viruses (as compared to other viral species) due to the rapid growth of cancer tissue. Indeed, as we have seen, several authors describing oncolytic viral dynamics at the whole-tumour scale have shown that the growth of the tumour is a significant obstacle to effective treatment (Wein et al. [27], Al-Johani et al. [184], Simbabwa et al. [185], for example). However, only one model at the cell scale has accounted for growth of the tissue: Jenner et al. considered a time–varying Voronoi lattice model for tumour cells, which also enabled the explicit representation of tumour cell proliferation and remodelling, and moreover facilitated the simulation of time–varying adjacency between cells [165]. Further work is needed, however, to more deeply explore the dynamics of infection on not just a realistic static cell packing, but also an evolving one.
@@ 5.3 Respiratory viruses
Figure 12: Spatially–structured models for influenza and SARS–CoV–2. Left: Average number of models for respiratory viruses per year. Centre: Respiratory virus models by model type. Right: Highly–cited publications for respiratory virus models over time. Highly–cited publications are as follows: Beauchemin et al. [29], Mitchell et al. [34], Getz et al. [9], Quirouette et al. [43], and Sego et al. [8].
@@
Viruses which infect the respiratory tract usually target the monolayer of epithelial cells which line the airways [159, 119, 15]. Thus, to an extent, the natural geometry of respiratory infection is two–dimensional, especially at the microscopic scale, and as a consequence virtually all of the respiratory infection models retained for analysis from our search also consider two spatial dimensions. However, at a larger length scale, there is substantial structural heterogeneity along the length of the respiratory tract. The respiratory tract has a complex, tree–like structure, comprising 24 generations of branching airways, such that upper passages like the trachea are relatively wide (on the order of centimetres), while deeper airways are very narrow (on the order of fractions of millimetres) [189, 159]. Ayadi et al. [190] and Moses et al. [119] have explicitly considered the dynamics of viral infection across the entire, heterogeneous lung. In the latter case, Moses et al. constructed a full–size, three–dimensional digital human lung, complete with a detailed branching structure for the cells [119]. The authors showed that viral dynamics on this highly complex structure were significantly different to the dynamics of the same infection model on a purely two–dimensional cell grid [119]. While the use of this hyperrealistic geometry demonstrates the notable differences in the dynamics of infections on flat sheets compared to the lung, there is a substantial conceptual leap between the two geometries. Other authors have instead employed an intermediate approach, studying the spread of viral infections on sections of the respiratory tract in order to identify specific ways in which the structure and geometry of the respiratory tract influence the spread of infection. Chen et al. [159] and Williams et al. [191] studied the spread of viral infections on individual airway branches, or on small sections of branching tissue. The latter group showed that tubular and branching tissue geometry imposes rich, spatially–directional infection dynamics [191]. These dynamics moreover exhibited an increased sensitivity to the rate of viral diffusion compared to the dynamics of infection on a flat cell sheet [191]. While these theoretical works have offered hints into the ways in which the dynamics of respiratory viral infections in vivo may differ from infections in vitro, a deeper integration with experimental data is now necessary to gain quantitative insights into the spread of infection in the respiratory tract. Obtaining relevant in vivo data remains a challenge, however, advances in imaging techniques and genetic engineering (discussed in Section 7.2) may soon provide data necessary for such analysis.
@@
Another characteristic feature of the lung is the mucociliary escalator, the directed flow of the mucus lining the respiratory epithelium from lower to upper regions of the respiratory tract. Quirouette et al. [43] included the mucociliary escalator in a model of influenza infection, and suggested that infection can effectively only spread up the respiratory tract from the initial point of virion deposition [43]. The rate of mucociliary flow is variable at different depths of the respiratory tract, with slower flow in lower airways [189, 159]. Chen et al. compared the dynamics of SARS–CoV–2 infection in airway tissue from different generations of the respiratory tract and suggested that infection spread was severely limited in the lower airways [159]. The authors postulated that infection propagation in the deep lung must therefore be driven by a large quantity of viral depositions from upper regions of the lung, such as virus–laden droplets or aerosols [159]. Chakravarty et al. — who extended the Quirouette model — reached a similar conclusion, arguing that aerosolisation of infected nasopharyngeal mucosa (ANM) was necessary for infection to reach the deep lung [192]. Vimalajeewa et al., by contrast, modelled advection down the respiratory tract in the course of SARS–CoV–2 infection due to air flow in the respiratory lumen (the authors accounted for the mucociliary escalator as a drag force and slip correction factor in the flow model) [55]. In addition to its role in disease progression, Zhang et al. suggested that the rate of mucus advection also influences the nasal viral load [163]. The authors argue that variation in this rate between individuals could offer one potential explanation for the extreme variation observed in clinical nasal swab titres among COVID–19 patients [163].
@@ 5.4 Other viruses
Figure 13: Spatially–structured models for other specific viruses. Left: Average number of models for other specific viruses per year. Centre: Other specific virus models by model type. Right: Highly–cited publications for other virus models over time. Highly–cited publications are as follows: Xiao et al. [31], Heilmann et al. (2010) [33], Heilmann et al. (2012) [36], Schiffer et al. [37], Eriksen et al. [41], and Simmons et al. [44].
@@
In addition to the viruses which we have discussed above, several other species have been considered. Among these, a substantive body of work has been developed for the modelling of bacteriophage (or simply “phage”) dynamics. Phage and the bacteria which they infect are both some of the largest biological populations in existence, and their interaction is one of the oldest predator–prey relationships on Earth [44, 193]. Bacteria and phage inhabit and interact within a broad array of environments, including in soil, water, and within larger organisms [33, 152]. There is a major difference between phage dynamics and the dynamics of the other virus species which we have considered in this review, in that while we have so far mainly considered viruses which infect tissues within a single host, the bacteria which phage infect are single–cell organisms. However, there are also many similarities. Phage–bacteria systems are frequently studied — and modelled — in vitro (Valdez et al. [194], de Rioja et al. [50], for example), and the dynamics of these systems can be directly compared to the in vitro dynamics of other, human–adapted viruses. Bacteria also frequently form dense, surface–bound microbial communities called biofilms, which can be likened to the crowded environment within tumours [194, 44, 41]. Despite first appearances, modelling of phage–bacteria dynamics is therefore closely interlinked with models for other viral species, and forms an important component of the development of spatially–structured viral dynamics models in general.
@@
Due to the unique properties of phage–bacteria dynamics, as compared to the dynamics of other viral species we have discussed, modellers have considered the spatially–structured interactions between phage and microbe communities on very large spatiotemporal scales. Garrido Zornoza et al. modelled the spread of phage among a microbial community at an ecological spatial scale (the only such model we found in our search), which described spatially explicit microbial habitats, from which both viruses and host microbes were subject to stochastic aerosolisation and advection with wind [152]. The authors showed that this influx–outflux process among habitats lead to the spatial spread of microbial communities but also resulted in destabilisation and occasional local extinctions in individual habitats [152]. While phage–bacteria systems exhibit spatially–structured dynamics on this ecological spatial scale, most authors instead model the dynamics of phage–bacteria interactions on far smaller length scales, particularly in biofilms. The crowded nature of these environments requires particular consideration by modellers. Martínez-Calvo et al., for instance, considered density-dependent motility of bacteria and showed that the spread of phage infection among the colony can disrupt bacterial aggregation [195]. Moreover, as with oncolytic viral infections in tumours, the densely–packed environment of biofilms almost entirely suppress viral diffusion, resulting in highly localised viral spread. Taylor et al., for example, showed that the high cell density found in biofilms promotes the formation of spatial clustering of infection [196]. Valdez et al. showed that, as a consequence of this highly localised spread, differently–shaped initial distributions of phage among bacterial communities gave rise to spatially distinct colony structures [194]. Modellers have accounted for the crowded nature of biofilms in various ways. Some, including Heilmann et al. (2012) [36] and Mitarai et al. [193], have implemented linear diffusion to describe the dispersal of cell–free phage with a very small diffusion coefficient compared to say, the Stokes–Einstein estimate (Akpinar et al. [98], Beauchemin et al. [168], Sego et al. [8], for example, took similar approaches for other viral species). Other authors have instead opted for non–linear forms of the diffusion term: Hunter et al., for example, considered density–dependent diffusion [197]; Simmons et al. explicitly considered physical blocking of phage motion by bacteria [44]. A direct comparison of these different potential mechanisms is lacking, but could be compared to experimental evidence to provide insight into the characteristics of phage dispersal in the crowded biofilm environment, which may also have consequences for other viral species, especially oncolytic viruses.
@@
Hepatitis B and C viruses, which cause chronic infections of the liver, have also been modelled by several authors. Despite the recent development of highly effective direct–acting antivirals for the latter, hepatitis viruses are still globally widespread with relatively high rates of mortality [201, 202]. Xiao et al. [31] and Cangelosi et al. [153] have argued that spatial heterogeneity of infection may play a role in immune evasion, supporting viral persistence in the liver. Goyal and Murray [99] and Blahut et al. [107] have also argued that direct cell–to–cell transmission of hepatitis viruses also contributes to the chronicity of infection (we discuss modelling for cell–to–cell infection in greater depth in Section 6). Moreover, within–cell dynamics appears to be key in hepatitis B dynamics. Cangelosi et al. showed that viral cccDNA represents one of the most robust obstacles to hepatitis B treatment, and suggested that even a single cccDNA molecule would be sufficient to reinfect the entire liver [153]. Interestingly, although the non–spatial modelling literature for hepatitis B and C is fairly extensive (reviewed by Dahari et al. [202] and Canini and Perelson [203], for example), there are relatively few (seven) articles which account for spatial structure. This is despite clear biological evidence of the role of spatial dynamics in the spread of hepatitis B and C infection, such as the presence and suspected high rate of cell–to–cell transmission [204, 99, 107], and experimentally observed clusters of infected cells in hepatitis C–infected livers [203, 205]. There is therefore substantial scope for future model development for hepatitis viruses which takes into account spatial effects in the dynamics.